# KEK-TH-2773 Quantum vacuum energy and geometry of extra dimension

### v1

@@ D.2.1 Basis functions
$\left[\partial_{y}^{2}+\left(\frac{n^{\prime}}{n}-\frac{b^{\prime}}{b}\right)\partial_{y}-\frac{b^{2}}{n^{2}}\rho^{2}\pm\frac{n^{\prime}}{n}bM-b^{2}M^{2}\right]f_{u}^{(\pm)}(y;\rho)$ $=0.$ (D.17)
@@
$f_{u}^{(\pm)}(y;\rho)$ $=e^{\rho{\cal U}_{u}(y)}\left[\frac{1}{2}+\frac{h_{u,1}^{(\pm)}(y)}{\rho}+\frac{h_{u,2}^{(\pm)}(y)}{\rho^{2}}+\frac{h_{u,3}^{(\pm)}(y)}{\rho^{3}}+\cdots\right],$ (D.18)
@@
$\hskip 42.67912pt+\left\{\pm Mn^{\prime}(y)-n(y)b(y)M^{2}\right\}h_{u,l-1}^{(\pm)}(y)\bigg].\;\;\;\;\;(l=2,3,4,\cdots)$ (D.19)
Considering the conditions in (B.7) and (B.9), the basis function $C_{u}^{(\pm)}(y;\rho^{2})$ is expressed as
$C_{u}^{(\pm)}(y;\rho^{2})$ $=f_{u}^{C(\pm)}(y;\rho)+f_{u}^{C(\pm)}(y;-\rho),$ (D.20)
@@
$f_{u}^{C(\pm)}(y;\rho)$ $=e^{\rho{\cal U}_{u}(y)}\left[\frac{1}{2}+\frac{h_{u,1}^{C(\pm)}(y)}{\rho}+\frac{h_{u,2}^{C(\pm)}(y)}{\rho^{2}}+\frac{h_{u,3}^{C(\pm)}(y)}{\rho^{3}}+\cdots\right],$ (D.21)
@@
$h_{u,2l}^{C(\pm)}(u)$ $=0,\;\;\;\;\;h_{u,2l+1}^{C(\pm)}(u)=-\frac{n(u)}{b(u)}h_{u,2l}^{C(\pm)\prime}(u).\;\;\;\;\;\left(l=1,2,3,\cdots\right)$ (D.22)
@@
$S_{u}^{(\pm)}(y;\rho^{2})$ $=\frac{n(u)}{\rho b(u)}\left[f_{u}^{S(\pm)}(y;\rho)-f_{u}^{S(\pm)}(y;-\rho)\right],$ (D.23)
@@
$f_{u}^{S(\pm)}(y;\rho)$ $=e^{\rho{\cal U}_{u}(y)}\left[\frac{1}{2}+\frac{h_{u,1}^{S(\pm)}(y)}{\rho}+\frac{h_{u,2}^{S(\pm)}(y)}{\rho^{2}}+\frac{h_{u,3}^{S(\pm)}(y)}{\rho^{3}}+\cdots\right],$ (D.24)
@@
$h_{u,2l-1}^{S(\pm)}(u)$ $=0,\;\;\;\;\;h_{u,2l}^{S(\pm)}(u)=-\frac{n(u)}{b(u)}h_{u,2l-1}^{S(\pm)\prime}(u).\;\;\;\;\;\left(l=1,2,3,\cdots\right)$ (D.25)
D.2.2 5D propagator
@@
$H_{L\lambda}(y;\rho^{2})$ $=e^{-\rho{\cal U}_{L}(y)}\left[\frac{1}{2}-\frac{\tilde{h}_{L,1}^{(+)}(y)}{\rho}+\frac{\tilde{h}_{L,2}^{(+)}(y)}{\rho^{2}}-\frac{\tilde{h}_{L,3}^{(+)}(y)}{\rho^{3}}+\cdots\right],$ (D.26)
@@
$\tilde{h}_{L,m}^{(+)}(y)$ $\equiv h_{L,m}^{C(+)}(y)-n(L)Mh_{L,m-1}^{S(+)}(y),\;\;\;\;\;\left(m=1,2,3,\cdots\right)$ (D.27)
@@
$\bar{G}_{\lambda\lambda<}(\rho^{2},y,y^{\prime})$ $=-\frac{e^{-\rho\Delta(y,y^{\prime})}}{2a_{\rm sc}^{3}\rho}\left[1+\frac{2{\cal K}_{1}^{\lambda\lambda}(y,y^{\prime})}{\rho}+\frac{2{\cal K}_{2}^{\lambda\lambda}(y,y^{\prime})}{\rho^{2}}+\frac{2{\cal K}_{3}^{\lambda\lambda}(y,y^{\prime})}{\rho^{3}}+\cdots\right],$ (D.28)
@@
$\quad-2\sum_{l=1}^{m-1}(-1)^{l}h_{L,l}^{S(-)}(0){\cal K}_{m-l}^{\chi\chi}(y,y^{\prime}),\;\;\;\;\;\left(m\geq 2\right)$ (D.29)
@@
${\cal K}_{m}^{\lambda\lambda}(y,y^{\prime})$ $\equiv\left.{\cal K}_{m}^{\chi\chi}(y,y^{\prime})\right|_{h_{0,l}^{S(-)}(y)\to\tilde{h}_{0,l}^{(-)}(y),\,h_{L,l}^{S(-)}(y^{\prime})\to\tilde{h}_{L,l}^{(-)}(y^{\prime})}.\;\;\;\;\;\left(m=1,2,3,\cdots\right)$ (D.30)
D.3 Expansion of other quantities
When $y^{\prime}=y+1/\Lambda_{\rm cut}$, $\Delta(y,y^{\prime})$ defined in (4.9) (or (D.13)) can be expanded as
$\Delta(y,y^{\prime})$ $=\frac{b(y)}{n(y)\Lambda_{\rm cut}}\left[1+\frac{X^{(1)}(y)}{\Lambda_{\rm cut}}+\frac{X^{(2)}(y)}{\Lambda_{\rm cut}^{2}}+\frac{X^{(3)}(y)}{\Lambda_{\rm cut}^{3}}+\cdots\right],$ (D.31)
@@ D.3 Expansion of other quantities
$X^{(l)}(y)$ $=\frac{n(y)}{(l+1)!b(y)}\partial_{y}^{l}\left(\frac{b(y)}{n(y)}\right).$ (D.32)
@@
$\Gamma(a,\delta)$ $=\Gamma(a)\left[1-\delta^{a}e^{-\delta}\left(\frac{1}{\Gamma(a+1)}+\frac{\delta}{\Gamma(a+2)}+\frac{\delta^{2}}{\Gamma(a+3)}+\cdots\right)\right],$ (D.34)
@@
$\quad-\frac{\Gamma(a)\bar{\rho}^{a}}{\Gamma(a+1)}\left[1+{\cal O}\left(\frac{1}{\Lambda_{\rm cut}}\right)\right],$ (D.35)

### v2

@@ E.2.1 Basis functions
$\left[\partial_{y}^{2}+\left(\frac{n^{\prime}}{n}-\frac{b^{\prime}}{b}\right)\partial_{y}-\frac{b^{2}}{n^{2}}\rho^{2}\pm\frac{n^{\prime}}{n}bM-b^{2}M^{2}\right]f_{u}^{(\pm)}(y;\rho)$ $=0.$ (E.17)
@@
$f_{u}^{(\pm)}(y;\rho)$ $=e^{\rho{\cal U}_{u}(y)}\left[\frac{1}{2}+\frac{h_{u,1}^{(\pm)}(y)}{\rho}+\frac{h_{u,2}^{(\pm)}(y)}{\rho^{2}}+\frac{h_{u,3}^{(\pm)}(y)}{\rho^{3}}+\cdots\right],$ (E.18)
@@
$\hskip 42.67912pt+\left\{\pm Mn^{\prime}(y)-n(y)b(y)M^{2}\right\}h_{u,l-1}^{(\pm)}(y)\bigg].\;\;\;\;\;(l=2,3,4,\cdots)$ (E.19)
Considering the conditions in (C.7) and (C.9), the basis function $C_{u}^{(\pm)}(y;\rho^{2})$ is expressed as
$C_{u}^{(\pm)}(y;\rho^{2})$ $=f_{u}^{C(\pm)}(y;\rho)+f_{u}^{C(\pm)}(y;-\rho),$ (E.20)
@@
$f_{u}^{C(\pm)}(y;\rho)$ $=e^{\rho{\cal U}_{u}(y)}\left[\frac{1}{2}+\frac{h_{u,1}^{C(\pm)}(y)}{\rho}+\frac{h_{u,2}^{C(\pm)}(y)}{\rho^{2}}+\frac{h_{u,3}^{C(\pm)}(y)}{\rho^{3}}+\cdots\right],$ (E.21)
@@
$h_{u,2l}^{C(\pm)}(u)$ $=0,\;\;\;\;\;h_{u,2l+1}^{C(\pm)}(u)=-\frac{n(u)}{b(u)}h_{u,2l}^{C(\pm)\prime}(u).\;\;\;\;\;\left(l=1,2,3,\cdots\right)$ (E.22)
@@
$S_{u}^{(\pm)}(y;\rho^{2})$ $=\frac{n(u)}{\rho b(u)}\left[f_{u}^{S(\pm)}(y;\rho)-f_{u}^{S(\pm)}(y;-\rho)\right],$ (E.23)
@@
$f_{u}^{S(\pm)}(y;\rho)$ $=e^{\rho{\cal U}_{u}(y)}\left[\frac{1}{2}+\frac{h_{u,1}^{S(\pm)}(y)}{\rho}+\frac{h_{u,2}^{S(\pm)}(y)}{\rho^{2}}+\frac{h_{u,3}^{S(\pm)}(y)}{\rho^{3}}+\cdots\right],$ (E.24)
@@
$h_{u,2l-1}^{S(\pm)}(u)$ $=0,\;\;\;\;\;h_{u,2l}^{S(\pm)}(u)=-\frac{n(u)}{b(u)}h_{u,2l-1}^{S(\pm)\prime}(u).\;\;\;\;\;\left(l=1,2,3,\cdots\right)$ (E.25)
E.2.2 5D propagator
@@
$H_{L\lambda}(y;\rho^{2})$ $=e^{-\rho{\cal U}_{L}(y)}\left[\frac{1}{2}-\frac{\tilde{h}_{L,1}^{(+)}(y)}{\rho}+\frac{\tilde{h}_{L,2}^{(+)}(y)}{\rho^{2}}-\frac{\tilde{h}_{L,3}^{(+)}(y)}{\rho^{3}}+\cdots\right],$ (E.26)
@@
$\tilde{h}_{L,m}^{(+)}(y)$ $\equiv h_{L,m}^{C(+)}(y)-n(L)Mh_{L,m-1}^{S(+)}(y),\;\;\;\;\;\left(m=1,2,3,\cdots\right)$ (E.27)
@@
$\bar{G}_{\lambda\lambda<}(\rho^{2},y,y^{\prime})$ $=-\frac{e^{-\rho\Delta(y,y^{\prime})}}{2a_{\rm sc}^{3}\rho}\left[1+\frac{2{\cal K}_{1}^{\lambda\lambda}(y,y^{\prime})}{\rho}+\frac{2{\cal K}_{2}^{\lambda\lambda}(y,y^{\prime})}{\rho^{2}}+\frac{2{\cal K}_{3}^{\lambda\lambda}(y,y^{\prime})}{\rho^{3}}+\cdots\right],$ (E.28)
@@
$\quad-2\sum_{l=1}^{m-1}(-1)^{l}h_{L,l}^{S(-)}(0){\cal K}_{m-l}^{\chi\chi}(y,y^{\prime}),\;\;\;\;\;\left(m\geq 2\right)$ (E.29)
@@
${\cal K}_{m}^{\lambda\lambda}(y,y^{\prime})$ $\equiv\left.{\cal K}_{m}^{\chi\chi}(y,y^{\prime})\right|_{h_{0,l}^{S(-)}(y)\to\tilde{h}_{0,l}^{(-)}(y),\,h_{L,l}^{S(-)}(y^{\prime})\to\tilde{h}_{L,l}^{(-)}(y^{\prime})}.\;\;\;\;\;\left(m=1,2,3,\cdots\right)$ (E.30)
E.3 Expansion of $\Delta(y,y^{\prime})$ and ${\cal W}(a,y)$
When $y^{\prime}=y+1/\Lambda_{\rm cut}$, $\Delta(y,y^{\prime})$ defined in (4.9) (or (E.13)) can be expanded as
$\Delta(y,y^{\prime})$ $=\frac{b(y)}{n(y)\Lambda_{\rm cut}}\left[1+\frac{X^{(1)}(y)}{\Lambda_{\rm cut}}+\frac{X^{(2)}(y)}{\Lambda_{\rm cut}^{2}}+\frac{X^{(3)}(y)}{\Lambda_{\rm cut}^{3}}+\cdots\right],$ (E.31)
@@ E.3 Expansion of $\Delta(y,y^{\prime})$ and ${\cal W}(a,y)$
$X^{(l)}(y)$ $=\frac{n(y)}{(l+1)!b(y)}\partial_{y}^{l}\left(\frac{b(y)}{n(y)}\right).$ (E.32)
@@
$\Gamma(a,\delta)$ $=\Gamma(a)\left[1-\delta^{a}e^{-\delta}\left(\frac{1}{\Gamma(a+1)}+\frac{\delta}{\Gamma(a+2)}+\frac{\delta^{2}}{\Gamma(a+3)}+\cdots\right)\right],$ (E.34)
@@
$\quad-\frac{\Gamma(a)\bar{\rho}^{a}}{\Gamma(a+1)}\left[1+{\cal O}\left(\frac{1}{\Lambda_{\rm cut}}\right)\right],$ (E.35)